On a roulette wheel, there are 38 slots, 18 of which are colored black. If you bet $5 on black and win, you get $10 back. If black doesn’t come up, you lose your $5. Find the expected value of playing the game 100 times.
Answer #1

From the given data, the following Table iscalculated:

Event Value (x) Probility (p) x p
Black comes up 10 - 5 = 5 18/38 = 0.4737 5 X 0.4737 = 2.3684
Black does not comes up - 5 20/38 = 0.5263 - 5 X 0.5263 = - 2.6316
Expected Value 2.3684 - 2.6316 = - 0.2632

So,

Expected value of playing the game 1 time =- 0.2632 (loss)

So,

Expected value of playing the game 100 times.= - 0.2632 X 100 = - 26.32 (loss)

So,

Answer is:

- 26.32 (loss)